A town situated 20km away from a power plant generating power at 440V , requires 600 Kw of electric power at 200 V . The resistance of two line wires carrying  power is 0.4 ohm per km . the towns gets power  from line through 3000-220 V step up transformer at a substation in the town


 


1) find the power losses in the form of heat?


2) how much power must the plant supply assuming there is a negligible power loss due to leakage ?


3) what is the voltage drop on the line ?


 


ans : 1) 640 KW


2) 1240 kW


3) 3200V

Grade:11

1 Answers

jyoti dogra
11 years ago

Answer :

 

Total resistance 0.4*40 = 16Ω

 

Irms on line = Power / Volt = ( 600 * 1000 ) / 3000 = 200 A

 

Line power loss = (Irms)2  * R = (200)2 * 16 = 640000 W = 640 KW

 

Power suplied = 600 + 640 = 1240 KW

 

Voltage drop = Irms * R = 200 * 16 = 3200 V

 

 

 

................. THANK YOU ......................

 

to all who didnt posted any reply to this so far as that made me to find a solution to this myself :)